A boy who is 4 feet 5 inches tall stands near a street light that is 15 feet tall. Find a function that models the length L (in feet) of his shadow in terms of his distance d feet from the base of the street light
Answer by scott8148(6628)

this is a problem using similar triangles

15/4.5=(d+L)/L ... 15L=4.5d+4.5L ... 10.5L=4.5d ... L=4.5d/10.5=3d/7